# Liquid water under vibrational strong coupling: an extended cavity Born-Oppenheimer molecular dynamics study

### v3

@@ Abstract
A computational study of liquid water when the system is coupled with a (model) Fabry-Perot cavity is reported. At this end, the Cavity Born-Oppenheimer Molecular Dynamics approach proposed recently (Li et al., Proc. Nat. Acad. Sci. USA, 2020, 117, 18324–18331) is employed and different properties of water under vibrational strong coupling (VSC) are investigated. Different cavity frequencies were considered, corresponding to different modes in the IR spectrum of liquid water: high frequency (corresponding to O–H stretching modes), medium frequency (corresponding to water molecule bending) and low frequencies (corresponding to librational modes). Simulations were done both using classical and quantum nuclear dynamics, this last with Ring Polymer Molecular Dynamics methods. Similar effect of the cavity are obtained in both cases. Namely, whereas the splitting of the infrared spectrum is obtained for all cavity frequencies, no effect on structural properties are obtained. Also transport and dynamical properties, including the diffusion coefficient, hydrogen bond (HB) relaxation and HB jump exchange times show no effect when coupled to the cavity and when an extended statistical analysis is performed. This lack of effect may be due to the model’s inability to account for bond breaking or the absence of molecular polarizability.
@@ II.1 Molecular Dynamics Simulations
We simulated a system composed by a box of 216 water molecules using the flexible q-TIP4P/F force field. Habershon, Markland, and Manolopoulos (2009) The box size was set at 18.64 Åto reproduce the experimental water density at 300 K, corresponding to what used by Li et al. Li, Subotnik, and Nitzan (2020) Standard periodic boundary conditions with an Ewald summation for the intermolecular Coulombic interactions were used. In the classical nuclear dynamics simulations, for each cavity frequency (and for the simulation without the cavity), we first equilibrated the system in the canonical ensemble at 300 K, using a Langevin thermostat with a friction caracteristic time = 100 fs, and a time step of 1 fs, using the velocity Verlet scheme. Verlet (1967) We then selected 30 initial conditions to run NVE simulations of 100 ps with a propagation time step of 0.5 fs to obtain all the properties except the dielectric constant where 20 NVE trajectories of 1 ns were performed and the IR and VDOS where 4 trajectories of 20 ps were carried out. In the RPMD simulations, the equilibration trajectories employed a white noise Langevin thermostat on the beads and a global velocity rescaling thermostat on the centroid mode, with = 100 fs. Ceriotti et al. (2010); Rossi, Ceriotti, and Manolopoulos (2014) For the dielectric constant, 50 T-RPMD trajectories of 300 ps were propagated in which only the beads are thermostatted, Ceriotti et al. (2010) using the positions and velocities extracted from the equilibrated trajectories, as initial conditions. For the dynamical properties, 30 NVE trajectories of 100 ps, also started from the equilibrated trajectories, were exploited when we used 8 beads and just two trajectories of 100 ps for 32 beads. All RPMD simulations employed a time step of 0.25 fs.
@@ II.2 Calculations of Properties
To determine the jump time, the cross-correlation function of the probability for a water OH group to donate a stable hydrogen-bond to an initial acceptor at time 0 and to form a hydrogen-bond to a new stable partner at time is used: Laage and Hynes (2008)
(7) This correlation function gives the probability to have formed a stable final state at time when the system was in the initial state at time . Thus the decay time of the complementary probability, , provides the jump time corresponding to the time for a water molecule to switch. Laage and Hynes (2008)
@@
Finally, it was shown that it is possible to estimate the free energy barrier corresponding to an hydrogen bond exchange from the oxygen-oxygen radial distribution function. Wilkins et al. (2017) In fact, this barrier corresponds to the free energy cost necessary to bring a couple of H-bonded water molecule and it is can be approximated with the free energy costs of elongation of the distance between the oxygen atom of leaving H-bond donor water and the initial acceptor molecule and the compression of the distance between the oxygen atom of the same H-bond donor molecule with the new acceptor molecule. The potential of mean force of this process, , is related to the oxygen-oxygen radial distribution function, :
@@ III.1 Polariton Formation
When the cavity frequency is at resonance with a water vibrational mode, a splitting can be seen in the resulting IR spectrum, corresponding to the formation of two polaritonic modes. The two new peaks correspond to the P+ and P- polaritonic modes that arise from the hybrid state as shown in Figure 1(b). On the other hand, for the VDOS spectra, shown in Figure 1(c), no splitting occurs, since no selection rules apply and so also the modes which are not hybridized with matter, the dark states (DS), are present. These states are numerically dominant over the bright states as depicted in Figure 1(a) and explained numerically in details by de la Fuente et al. de la Fuente, R.Spezia, and Vuilleumier (2024)
@@
In Figure 2 (panels A and B), we show in details the IR spectra obtained under resonance conditions: in particular we show both the spectrum calculating from dipoles in the (x,y) plane, corresponding to the plane on which the field is coupled to the matter, and in the z-axis. In panel A we show the spectra obtained when the cavity is set to 600 cm-1, while in panel B when it is set to 1650 cm-1. Interestingly, we notice that for the signals in the light polarization plane, where the polaritons are formed, the other bands, which are out-of-resonance with the cavity, are also shifted. When the cavity is in resonance with the librational motion we observe that the bending is blue shifted by about 80 cm-1 (the maximum being now at about 1730 cm-1) and the stretching is also blue shifted by about 100 cm-1 (the maximum is now at about 3650 cm-1). The same phenomenon is observed when the cavity is resonant with the bending: now the broad librational broad band is red-shifted, while the stretching is still blue shifted. This phenomenon is probably due to the fact that the cavity is off-resonance with the other modes, while it remains close enough to slightly shift them.
@@
To verify that this phenomenon, we have employed a model Pauli-Fierz Hamiltonian where the matter is represented by three modes corresponding to the frequencies of liquid water. To represent the broadening of liquid phase, a distribution of frequencies is considered for each mode. The cavity is then added at a frequency of 1650 cm -1. With this Hamiltonian, Path Integral Monte Carlo simulations allow to obtain IR spectrum and polaritons in very good agreement with exact quantum Hamiltonian, as shown recently by some of us. de la Fuente, R.Spezia, and Vuilleumier (2024) More details are reported in Appendix A.
@@
In panel C of Figure 2, we report results obtained with this 3-mode model with the cavity set at 1650 cm -1. Notably, we observe that the high frequency peak is blue shifted while the low frequency is red shifted. This confirms that the Born-Oppenheimer cavity-MD approach is able to correctly reproduce spectroscopic properties for systems under VSC, even concerning detailed features such as the shift of modes not directly at resonance with the cavity.
@@ III.2 Structural and Thermodynamic Properties
To continue the study of water within the cavity, we look at the dielectric constant, , obtained as function of cavity frequency. In this case, the simulation time of each trajectory is increased to 20 ns due the well-known slow convergence of this property. Results are summarized in Table 1 which also includes values reported in the literature using the same water model. Also in this case, we cannot see any significant effect on the cavity. The results found in the literature with the same water model (without the cavity or with the cavity set at 3550 cm-1) also correspond to our findings.
The fact that no effect is observed for structural and thermodynamic properties, while searched extensively, should be expected for classical particles under the Pauli-Fierz Hamiltonian, as a consequence of classical statistical mechanics. The reasoning is detailed in Appendix B.
@@ III.3 Transport and Dynamical Properties
52.7 6.5111From Eltareb et al. Eltareb, Lopez, and Giovambattista (2021) | ||||||
@@ III.4 Statistical Discussion
As pointed out previously, the evaluation of statistical uncertainty is a key aspect to decipher if a property can be considered modified or not. It is also important to evaluate how many trajectories are necessary to have a reasonable convergence. At this end we have employed the bootstrapping method, increasing the number of trajectories up to 60 in evaluating transport and dynamical properties (D, and ). Using bootstrapping, a dataset of size with is randomly selected from the extended dataset, and the mean of this sample of size is then calculated. For each , this random selection is repeated 1000 times from which we can determine an error reported here as the 95% confidence interval. As discussed previously, one could argue that for = 1650 cm -1 there can be some effect of the cavity: we have thus considered this cavity frequency in the bootstrapping analysis.
@@ III.5 Nuclear Quantum Effects
No Cavity | = 600 cm -1 | |
55 2 | 58 2 | |
53 7111From Eltareb et al. Eltareb, Lopez, and Giovambattista (2021) | ||
(103 Åps-1) | 246 6 | 247 7 |
220111From Eltareb et al. Eltareb, Lopez, and Giovambattista (2021) | ||
221222From Habershon et al. Habershon, Markland, and Manolopoulos (2009) 1 | ||
(ps) | 2.54 0.07 | 2.51 0.07 |
@@ IV Conclusions
Concerning the diffusion coefficient, experiments have shown that the diffusion of ions can be modified by VSC and in case of water this is likely due to a difference in proton conductivity. Fukushima, Yoshimitsu, and Murakoshi (2022, 2023) The non reactive q-TIP4P/F model clearly cannot provide any support on it and our study shows that with this popular water model, the diffusion coefficient is not altered by VSC. Unfortunately there are no experimental studies on the H-bond reorentational times under VSC. H-bond jump times are not accessible experimentally, but were studied here since they hold all the properties of a chemical reaction even if no covalent bond is broken/made, but only H-bonds. Also on those properties related to H-bond dynamics, using the present model we cannot highlight any effect of VSC at any cavity frequency.
As pointed out, we have extensively studied liquid water under VSC employing the q-TIP4P/F water model. This model has two limitations which could be at the origin of the absence of any relevant effect of VSC on dynamical properties: (i) it is a fixed charge model (the cavity field does not change the charge distribution of water molecules); (ii) the water model does not allow bond breaking/making. Nevertheless, the model is able to recover the spectroscopic features, not only in terms of formation of polaritonic bands in the IR signal but also the small blue and red shifts of other bands not in resonance with the cavity mode.
The fixed cavity model could be improved in the future considering molecular polarizability, as a first step. In fact, recent theoretical studies by Rubio and co-workers have shown that VSC could modify the polarizability of molecules present in the cavity. Sidler et al. (2024) Unfortunately, their method is based on very expensive calculations which cannot be presently applied to liquid systems. Water models including polarizability can be employed as a first approximated description of fluctuating molecular electronic density (via induced dipoles, for example) and the Pauli-Fierz Hamiltonian can be adapted. Additionally, an interesting question is whether the inclusion of polarisation, in classical MD, through models such as the fluctuating charges model, the Drude model or the induced dipoles model is sufficient or is the consideration of the electronic structure necessary. In any case, it will be important to both consider a number of molecules under VSC which is sufficiently large (at least 100) de la Fuente, R.Spezia, and Vuilleumier (2024) and have a statistical convergence of the properties. Our research are now going in that direction.

### v4

@@ Abstract
A computational study of liquid water when the system is coupled with a (model) Fabry-Perot cavity is reported. At this end, the Cavity Born-Oppenheimer Molecular Dynamics approach proposed recently (Li et al., Proc. Nat. Acad. Sci. USA, 2020, 117, 18324–18331) is employed and different properties of water under vibrational strong coupling (VSC) are investigated. Different cavity frequencies are considered, corresponding to different modes in the IR spectrum of liquid water: high frequency (corresponding to O–H stretching modes), medium frequency (corresponding to water molecule bending) and low frequencies (corresponding to librational modes). Simulations were done both using classical and quantum nuclear dynamics, this last via Ring Polymer Molecular Dynamics. Similar effects of the cavity are obtained in both cases. Namely, whereas the infrared spectrum is observed to be split for all cavity frequencies, no effects on structural properties are detected. In addition, transport and dynamical properties, including the diffusion coefficient, molecular reorientation and hydrogen bond (HB) jump exchange times, show no effect due to cavity coupling when an extended statistical analysis is performed.
@@ II.1 Molecular Dynamics Simulations
We simulated a system composed by a box of 216 water molecules using the flexible q-TIP4P/F force field. Habershon, Markland, and Manolopoulos (2009) The box size was set at 18.64 Å to reproduce the experimental water density at 300 K, corresponding to what used by Li et al. Li, Subotnik, and Nitzan (2020) Standard periodic boundary conditions with an Ewald summation for the intermolecular Coulombic interactions were used. In the classical nuclear dynamics simulations, for each cavity frequency (and for the simulation without the cavity), we first equilibrated the system in the canonical ensemble at 300 K, using a Langevin thermostat with a friction caracteristic time = 100 fs, and a time step of 1 fs, using the velocity Verlet scheme. Verlet (1967) We then selected 30 initial conditions to run NVE simulations of 100 ps with a propagation time step of 0.5 fs to obtain all the properties except the dielectric constant where 20 NVE trajectories of 1 ns were performed and the IR and VDOS where 4 trajectories of 20 ps were carried out. In the RPMD simulations, the equilibration trajectories employed a white noise Langevin thermostat on the beads and a global velocity rescaling thermostat on the centroid mode, with = 100 fs. Ceriotti et al. (2010); Rossi, Ceriotti, and Manolopoulos (2014) For the dielectric constant, we then extracted 50 initial conditions and ran T-RPMD trajectories of 300 ps each, in which only the internal modes are thermostatted, Ceriotti et al. (2010) using the positions and velocities extracted from the equilibrated trajectories as initial conditions. For the dynamical properties, 30 NVE trajectories of 100 ps, also started from the equilibrated trajectories, were exploited when we used 8 beads and just two trajectories of 100 ps for 32 beads. All RPMD simulations employed a time step of 0.25 fs.
@@
As pointed out by Li et al. Li, Subotnik, and Nitzan (2020), for lower cavity modes overtones of the cavity should also be included. We have then run additional trajectories setting the cavity fundamental frequency at 600 cm-1, and including additional cavity overtone modes at 1200, 1800, 2400, 3000 and 3600 cm-1 using the multi-cavity method described by Li et al. Li, Subotnik, and Nitzan (2020) and for two values of coupling strength of and a.u. .
@@ II.2 Calculations of Properties
To determine the jump time, the cross-correlation function of the probability for a water OH group to donate a stable hydrogen-bond to an initial acceptor at time 0, , and to form a hydrogen-bond to a new stable partner at time , , is used: Laage and Hynes (2008)
(7) Here, when the OH group under consideration donates a stable H-bond to at time and otherwise. The H-bond presence is determined from strict geometric criteria ( Å, Å, and )Laage and Hynes (2008). This correlation function gives the probability to form a stable final H-bonded state at time when the system was engaged in a stable H-bond with a different acceptor in the initial state at time . Absorbing boundary conditions in the product state are applied to consider the first acceptor exchange only. Thus the decay time of the complementary probability, , provides the jump time corresponding to the time for a water molecule to switch stable H-bond acceptors. Laage and Hynes (2008)
@@
Finally, it was shown that the free energy barrier corresponding to an hydrogen bond exchange can be estimated from the oxygen-oxygen radial distribution function. Wilkins et al. (2017) This free energy barrier is the sum of two terms, which correspond respectively to the elongation of the distance between the oxygen atom of the jumping H-bond donor water and the initial acceptor molecule, and to the compression of the distance between the oxygen atom of the same H-bond donor molecule and the new acceptor molecule. Each term can be determined from the potential of mean force along the oxygen-oxygen distance, , which is related to the oxygen-oxygen radial distribution function (RDF), :
@@ III.1 Polariton Formation
When the cavity frequency is at resonance with a water vibrational mode, a splitting can be seen in the resulting IR spectrum, corresponding to the formation of two polaritonic modes. The two new peaks correspond to the P+ and P- polaritonic modes that arise from the hybrid state as shown in the IR spectra reported in Figure 1(b). On the other hand, for the VDOS spectra, shown in Figure 1(c), no splitting occurs, since no selection rules apply and so also the modes which are not hybridized with matter, the dark states (DS), are present. These states are numerically dominant over the bright states as depicted in Figure 1(a) and explained numerically in details by de la Fuente et al. de la Fuente, R.Spezia, and Vuilleumier (2024)
@@
In Figure 2 (panels A and B), we show in details the IR spectra obtained under resonance conditions: in particular we show both the spectrum calculating from dipoles in the (x,y) plane, corresponding to the plane on which the field is coupled to the matter, and in the z-axis. In panel A, we show the spectra obtained when the cavity is set to 600 cm-1, while in panel B when it is set to 1650 cm-1. Interestingly, we notice that for the signals in the light polarization plane, where the polaritons are formed, the other bands, which are out-of-resonance with the cavity, are also shifted. When the cavity is in resonance with the librational motion, we observe that the bending is blue shifted by about 80 cm-1 (the maximum being now at about 1730 cm-1) and the stretching is also blue shifted by about 100 cm-1 (the maximum is now at about 3650 cm-1). The same phenomenon is observed when the cavity is resonant with the bending: now the broad librational broad band is red-shifted, while the stretching is still blue shifted. This phenomenon is probably due to the fact that the cavity is off-resonance with the other modes, while it remains close enough to slightly shift them. These spectral features, namely the asymmetry of the polaritons and the shifts of the other bands, are obtained also by employing a simple harmonic oscillator model proposed by Lieberherr et al. Lieberherr et al. (2023) in which the cavity-free IR spectrum was given as unperturbed input. Results are shown in the same Figure 2, panels A and B.
@@
To understand these particular features of polaritonic bands, we have also employed a model Pauli-Fierz Hamiltonian where the matter is represented by three modes corresponding to the frequencies of liquid water. To represent the broadening of liquid phase, a distribution of frequencies is considered for each mode, reflecting the distribution width obtained from cavity-free simulations. The cavity is then added at a frequency of 1650 cm -1. With this Hamiltonian, Path Integral Monte Carlo simulations allow to obtain IR spectrum and polaritons in very good agreement with exact quantum Hamiltonian, as shown recently by some of us. de la Fuente, R.Spezia, and Vuilleumier (2024) More details are reported in Appendix A.
@@
In panel C of Figure 2, we report results obtained with this 3-mode model with the cavity set at 1650 cm -1 (note that the non exactly unimodal IR bands of the uncoupled system is due to the numerical sampling of 100 vibrational modes). Notably, we observe that the high frequency peak is blue shifted while the low frequency is red shifted. This confirms that the Born-Oppenheimer cavity-MD approach is able to correctly reproduce spectroscopic properties for systems under VSC, even concerning detailed features such as the shift of modes not directly at resonance with the cavity.
We also obtain that the lower polariton (LP) is much less intense than the upper polariton (UP), as in molecular simulations. From PIMC simulations, we can obtain the coefficients of the cavity and of the matter modes in the coupled system which reflect the light-matter hybridization and can be used to get deeper details. In particular, we have analyzed how the cavity coefficients are distributed as a function of frequency. In the simple case of a "isolated" polariton, the cavity contribution is 50 % on the LP and 50% on the UP. de la Fuente, R.Spezia, and Vuilleumier (2024) Here, however, we observe that the cavity has a non negligible contribution also away from the polaritonic band. Notably we divided the spectrum in four regions: (I) < 1300 cm-1; (II) 1300 < < 1650 cm-1; (III) 1650 < < 2500 cm-1 and (IV) > 2500 cm-1. Regions II and III correspond to LP and UP. From the cavity coefficients obtained from PIMC simulations we got that while the UP (region III) has almost 50% of the cavity, the LP (region II) has about 35% of the cavity and region I has about 15%. This explains the decrease in intensity of the LP.
Furthermore, from PIMC simulations it is possible to evaluate the degree of delocalization of the different bands by calculating the Inverse Participation Ratio (IPR) :
where are the coefficients for each matter mode of the final band and are the number of matter modes. In our previous study where only one vibrational band was considered, we obtained that the polaritonic bands have very low values, corresponding to high delocalization (note that by construction the maximum delocalization is , so 0.01 in the present model). Here we obtain that the upper polariton has a IPRUP = 0.012, while the lower polariton has IPRLP = 0.16. This means that the upper polariton is a "normal" polariton (i.e. as obtained when studying a single isolated vibrational mode), such that its delocalization is close to the maximal one. On the other hand, the lower polariton is more localized, with a value similar to what we obtained for a dark state. de la Fuente, R.Spezia, and Vuilleumier (2024) It is still bright (but much less intense than UP) because the cavity coefficient is not zero even if lower than for UP because it is spread over the LP and lower frequencies (the librational band in the present case).
@@ III.2 Structural and Thermodynamic Properties
To continue the study of water within the cavity, we look at the dielectric constant, , obtained as function of cavity frequency. In this case, the total simulation time is increased from 3 ns to 20 ns due the well-known slow convergence of this property. Results are summarized in Table 1 which also includes values reported in the literature using the same water model. Also in this case, we cannot see any significant effect on the cavity. The results found in the literature with the same water model (without the cavity or with the cavity set at 3550 cm-1) also correspond to our findings.
The fact that no effect is observed for structural and thermodynamic properties, while searched extensively, should be expected for classical particles under the Pauli-Fierz Hamiltonian, as a consequence of classical statistical mechanics. A concise derivation is given by Li et al. Li, Subotnik, and Nitzan (2020) and in the Appendix B we provide a more extended derivation.
@@ III.3 Transport and Dynamical Properties
| 52.7 6.5111From Eltareb et al. Eltareb, Lopez, and Giovambattista (2021) | ||||||
@@
Finally, we have tested if the multi-cavity description could have an impact on such properties, with a fundamental frequency of the cavity at 600 cm-1. Notably, for the 6.8 10-5 a.u coupling strength we obtained D = 0.195 0.030 Å ps-1, = 3.26 0.23 ps and = 3.97 0.24 ps, while for the coupling strength of 2.0 10-4 a.u., we got D = 0.209 0.018 Å ps-1, = 3.04 0.24 ps and = 3.67 0.26 ps. These values are, within the uncertainty, equivalent to what obtained without the cavity or with a single cavity mode.
@@ III.4 Statistical Discussion
As pointed out previously, the evaluation of statistical uncertainty is a key aspect to decipher if a property can be considered modified or not. It is also important to evaluate how many trajectories are necessary to have a reasonable convergence. At this end, we have employed the bootstrapping method, increasing the number of trajectories up to 60 in evaluating transport and dynamical properties (D, and ). Using bootstrapping, a dataset of size with is randomly selected from the extended dataset, and the mean of this sample of size is then calculated. For each , this random selection is repeated 1000 times from which we can determine an error reported here as the 95% confidence interval. As discussed previously, one could argue that for = 1650 cm -1 there can be some effect of the cavity: we have thus considered this cavity frequency in the bootstrapping analysis.
@@ III.5 Nuclear Quantum Effects
| No Cavity | = 600 cm -1 | = 3400 cm -1 | |
| 55 2 | 58 2 | ||
| 53 7111From Eltareb et al. Eltareb, Lopez, and Giovambattista (2021) | |||
| 59333From Li et al. Li et al. (2022) | 56.6333From Li et al. Li et al. (2022) | ||
| (10-3 Åps-1) | 246 6 | 247 7 | |
| 220111From Eltareb et al. Eltareb, Lopez, and Giovambattista (2021) | |||
| 221222From Habershon et al. Habershon, Markland, and Manolopoulos (2009) 1 | |||
| (ps) | 2.54 0.07 | 2.51 0.07 | |
@@ IV Conclusions
Concerning the diffusion coefficient, experiments have shown that the diffusion of ions can be modified by VSC and in case of water this is likely due to a difference in proton conductivity. Fukushima, Yoshimitsu, and Murakoshi (2022, 2023) The non reactive q-TIP4P/F model clearly cannot provide any support on it and our study shows that with this popular water model, the diffusion coefficient is not altered by VSC. Unfortunately there are no experimental studies on the H-bond reorentational times under VSC. H-bond jump times are not accessible experimentally, but were studied here since they hold all the properties of a chemical reaction even if no covalent bond is broken/made, but only H-bonds. Also on those properties related to H-bond dynamics, using the present model, we cannot highlight any effect of VSC at any cavity frequency.
As pointed out, we have extensively studied liquid water under VSC employing the q-TIP4P/F water model. This model has two limitations which could be at the origin of the absence of any relevant effect of VSC on dynamical properties: (i) it is a fixed charge model (the cavity field does not change the charge distribution of water molecules); (ii) the water model does not allow bond breaking/making. Nevertheless, the model is able to recover the spectroscopic features, not only in terms of formation of polaritonic bands in the IR signal but also the small blue and red shifts of other bands not in resonance with the cavity mode. As pointed out by Lieberherr et al. Lieberherr et al. (2023), the coupling strength used to reproduce the typical Rabi splitting using the q-TIP4P/F water model is unphysically high. This provides additional support to the conclusions of our work : since an extremely large coupling strength does not lead to any significant impact on any of the investigated water properties, one can safely suggest that lower coupling values should not cause any meaningful change in the liquid water properties. In addition, we note that while the qTIP4P-F model already provides an excellent description of the properties of water, future extensions of this work will examine VSC effects with more sophisticated water models.
The fixed cavity model could be improved in the future considering molecular polarizability, as a first step. In fact, recent theoretical studies by Rubio and co-workers have shown that VSC could modify the polarizability of molecules present in the cavity. Sidler et al. (2024) Unfortunately, their method is based on very expensive calculations which cannot be presently applied to liquid systems. Water models including polarizability can be employed as a first approximated description of fluctuating molecular electronic density (via induced dipoles, for example) and the Pauli-Fierz Hamiltonian can be adapted. Additionally, an interesting question is whether the inclusion of polarization, in classical MD, through models such as the fluctuating charges model, the Drude model or the induced dipoles model is sufficient or is the consideration of the electronic structure necessary. In any case, it will be important to both consider a number of molecules under VSC which is sufficiently large (at least 100) de la Fuente, R.Spezia, and Vuilleumier (2024) and have a statistical convergence of the properties. Our research are now going in that direction.